# Can AI be enabled to dynamical downscaling? Training a Latent Diffusion Model to mimic km-scale COSMO-CLM downscaling of ERA5 over Italy

### v1

@@ 5.6 Run-time performance
In this section, we compare the run-time performance of our tested models. These characteristics are of fundamental importance given the potential target applications of such models. Table 2 reports data for each model: to give a whole picture of the needed resources we provide information on both the training and inference requirements. The hour budgets indicated for LDM_res include hours to train/run the UNET also, which is needed to provide the residual data. LDM_res budgets are therefore comprehensive of the required computational time to train and run the whole modeling chain from scratch. Simulations were run either on NVIDIA GeForce RTX 4090 or NVIDIA A100 GPUs. The training dataset comprises  129000 hour-samples over a target domain of 576x672 pixels (at high resolution) and 72x86 pixels (at low-resolution). The UNET and GAN training ran with batches of and of for 2-m temperature and 10-m wind speed components, respectively. LDM_res training ran with 16- and 8- size batches for the corresponding two target variables. Inference times are calculated running with single-dimension batches.
As shown in Table 2, LDM_res implies more expensive training and inference processes when compared with the tested DL baselines. This evidence is expected, given the more complex structure of the diffusion model and its dimensions in terms of trainable parameters, which is an order of magnitude greater than that of the baselines. Nevertheless, the required computational time for both training and inference remains contained and competitive with the other available options. LDM_res requires  10 days over 8 GPUs to train both models and  30 hours on a single GPU to downscale one year of hourly data. As a term of comparison, we note here that a 1-year-long COSMO-CLM simulation ran over the very same domain requires 61 h, running on 2160 cores [2] (of course producing many more high-resolution variables than the sole 2-m temperature and 10-m horizontal wind components).
@@ 6 Discussion and Conclusions
However, the great performance of LDM_res comes at the cost of significantly higher computational requirements for both training and inference, when compared to the other DL models, the UNET or the GAN. Nonetheless, LDM_res still offers a significant advantage in terms of inference speed and computational efficiency once the model is trained when compared to the extensive computational resources required by COSMO-CLM.
@@ 7 Future work
The results presented in this work suggest several promising directions for further investigation into the application of latent diffusion models for downscaling. Future research could explore the ensemble generation capabilities of our Latent Diffusion Model, its effectiveness in downscaling discontinuous and chaotic variables such as precipitation—crucial for many applications—and the temporal consistency of the downscaled data, including methods to enforce this consistency within the model architecture. Long-term developments may include applying latent diffusion models to real-time weather forecasts, seasonal forecasts, and climate projections, with adjustments in the training procedure, particularly in selecting low-resolution input predictors and target reference truths.
Acknowledgments
Disclaimer
This Work has not yet been peer-reviewed and is provided by the contributing Author(s) as a means to ensure timely dissemination of scholarly and technical Work on a noncommercial basis. Copyright and all rights therein are maintained by the Author(s) or by other copyright owners. It is understood that all persons copying this information will adhere to the terms and constraints invoked by each Author’s copyright. This Work may not be reposted without explicit permission of the copyright owner.
8 Additional Material
8.1 Additional random snapshots of downscaled data
Some additional snapshots of downscaled data from all the tested models are shown in the following.
Figure 11: Downscaled variables from all the tested models against low-resolution ERA5 input data and high-resolution COSMO-CLM reference truth, for a randomly picked timestamp. The left columns refer to 2-m temperature, and the right columns refer to 10-m wind speed. The second and fourth columns show a zoom-in on Sardinia Island.
8.2 Calculation of verification metrics
In Section 5.2, we discuss the following metrics: three distance metrics, the Root Mean Square Error (RMSE), the mean bias (BIAS), the coefficient of determination (R2), and one correlation metric, the Pearson Correlation Coefficient (PCC). All the metrics are calculated with the xskillscore library [49]. Definitions are the following:
@@ 8.2 Calculation of verification metrics
8.3 On the contribution of the VAE
Figure 12 compares spectra from the COSMO-CLM test reference truth data with spectra from high-resolution test data generated using three models: the quadratic interpolation of ERA5 data, VAE_res, and LDM_res. This figure aims to highlight the decompression stage’s contribution from the latent space to the pixel space in LDM_res.
VAE_res takes as input the original reference truth high-resolution target variables. LDM_res takes as input the low-resolution ERA5 predictors, the high-resolution static data, and random noise, produces information in the latent space, and projects them in the pixel space using the VAE Decoder. Both VAE_res and LDM_res also use the corresponding UNET estimates for each target variable, but these quantities are subtracted before the encoding step (when applied) and added back after the decompression stage, acting essentially as constants. Therefore, the reconstruction errors in the spectra for VAE_res are thus solely attributed to the compression/decompression processes, while the reconstruction errors in the spectra for LDM_res arise from both the diffusion and decompression processes.
For 2-m temperature, the spectra from the VAE_res and LDM_res are nearly identical across all wavelengths, including the smallest scales. This indicates that the errors in reconstructing the spectra with LDM_res are attributable solely to the decompression stage, with the diffusion process effectively and accurately extracting latent features from the conditional data. On the contrary, for 10-m wind speed, which is a more chaotic field, discrepancies between the spectra from LDM_res and VAE_res are observed at scales smaller than approximately 7 km. These differences highlight the errors introduced by the sole extraction of features by the diffusion process.
Figure 12: Comparison of Radially Averaged Power Spectral Density (RAPSD) distributions for results from LDM_res and the LDM VAE against COSMO-CLM reference-truth. The left column refers to the 2-m temperature, and the right column refers to the 10-m wind speed. The first row shows the averaged-in-time spectra, across the whole test dataset. Notice that in the first row y-axes are logarithmic to highlight the tail of the distributions, hence the high frequencies. The bottom rows show the distributions of single-time RAPSD values for fixed wavelengths, namely 269, 20, 9, and 5 km.
8.4 LDM residual versus non-residual results
In this section, we report the comparison of results from the LDM trained with and without the residual approach, to provide an overview of the improvements introduced by this approach. We compare results in terms of frequency distribution and of radially average power spectral density, shown in Figures 13 and 14.
@@ 8.4 LDM residual versus non-residual results
The corresponding VAEs for the two models (VAE and VAE_res) perform the same except for the small scales of the 2-m temperature power spectra (not shown). Thus, the lack of performance of LDM can be attributed to the VAE only for this occasion (i.e., (iii)). All the other listed deficiencies are to be attributed to the diffusion process itself, which, when trained to reconstruct a residual field instead of the original target variable field, gains performance in the reconstruction of all the frequency distributions, and of the power spectra across all wavelengths (not only the small scales), especially for chaotic variables such as wind speed.

### v2

@@ 4.6 Run-time performance
In this section, we compare the run-time performance of our tested models. These characteristics are of fundamental importance given the potential target applications of such models. Table 2 reports data for each model: to give a whole picture of the needed resources we provide information on both the training and inference requirements. The hour budgets indicated for LDM_res include hours to train/run the UNET, which is needed to provide the residual data, and the VAE_res. LDM_res budgets are therefore comprehensive of the required computational time to train and run the whole modeling chain from scratch. Simulations were run either on NVIDIA GeForce RTX 4090 or NVIDIA A100 GPUs. The training dataset comprises  129000 hourly samples over a target domain of 576x672 pixels (at high resolution) and 72x86 pixels (at low resolution). The UNET and GAN training ran with 4-size batches for both target variables on the whole target domain. LDM_res training ran with 8- and 4-size batches for the two target variables (2-m temperature and 10-m wind components, respectively) on patches of 512x512 and 64x64 for high-resolution and low-resolution data, respectively. Inference times are calculated running with single-dimension batches, on the size of the entire domain.
As shown in Table 2, LDM_res implies more expensive training and inference processes when compared with the tested DL baselines. This evidence is expected, given the more complex structure of the diffusion model and its dimensions in terms of trainable parameters, which is an order of magnitude greater than that of the baselines. Nevertheless, the required computational time for both training and inference remains contained and competitive with the other available options. LDM_res requires  10 days over 8 GPUs to train the models for both the target variables and  30 hours on a single GPU to downscale one year of hourly temperature and wind data. As a term of comparison, we note here that a 1-year-long COSMO_CLM simulation ran over the very same domain requires 61 h, running on 2160 cores (Raffa et al., 2021) (of course producing many more high-resolution variables than the sole 2-m temperature and 10-m horizontal wind components).
@@ 5 Discussion and Conclusions
However, the great performance of LDM_res comes at the cost of significantly higher computational requirements for both training and inference, when compared to the other DL models, the UNET or the GAN. Nonetheless, LDM_res still offers a significant advantage in terms of inference speed and computational efficiency once the model is trained when compared to the extensive computational resources required by COSMO_CLM.
@@ 6 Future work
The results presented in this work suggest several promising directions for further investigation into the application of latent diffusion models for downscaling. Future research could explore the ensemble generation capabilities of our Latent Diffusion Model, its effectiveness in downscaling discontinuous and chaotic variables such as precipitation, crucial for many applications, and the temporal consistency of the downscaled data, including methods to enforce this consistency within the model architecture. Long-term developments may include applying latent diffusion models to real-time weather forecasts, seasonal forecasts, and climate projections, with adjustments in the training procedure, particularly in selecting low-resolution input predictors and target reference truths.
[…]
Data from ERA5 (Hersbach et al., 2020) was used as input low-resolution data for our models and was downloaded from the Copernicus Climate Change Service (Hersbach et al., 2023a, b). The results contain modified Copernicus Climate Change Service information 2020. Neither the European Commission nor ECMWF is responsible for any use that may be made of the Copernicus information or data it contains. The target high-resolution VHR-REA_IT data (Raffa et al., 2021; Adinolfi et al., 2023) was downloaded from the Euro-Mediterranean Center on Climate Change (CMCC) Data Delivery System (CMCC, 2021). CMCC produced the VHR-REA_IT dataset as a part of the Highlander project and released it under the CC BY 4.0 Licence. The full, preprocessed dataset used for the presented experiments is available on Zenodo (Tomasi et al., 2024c, d, e, f, g, h, i). Additionally, Zenodo hosts a sample dataset (Tomasi et al., 2024a) to test and train the models and the trained models themselves (Tomasi et al., 2024b). A dedicated GitHub repository (https://github.com/DSIP-FBK/DiffScaler) hosts the Pytorch Lightning (Falcon and The PyTorch Lightning team, 2019) code of the models described in this paper, based on the Lightning-Hydra-Template (Yadan, 2019), licensed under the MIT License. The repository also hosts the code to reproduce all the images shown in this paper. LDM_res v1.0 GitHub release is archived on Zenodo (Franch et al., 2024) and allows to download the code to reproduce the presented experiments.
Appendix A Additional snapshots of downscaled data
Some additional snapshots of downscaled data from all the tested models are shown in the following Figures 11 and 12. The left columns refer to 2-m temperature, and the right columns refer to 10-m wind speed. The second and fourth columns show a zoom-in on Sardinia Island.
Figure 11: Downscaled variables from all the tested models against low-resolution ERA5 input data and high-resolution COSMO-CLM reference truth, for an additional random timestamp.
Figure 12: Downscaled variables from all the tested models against low-resolution ERA5 input data and high-resolution COSMO-CLM reference truth, for an additional random timestamp.
Appendix B Calculation of verification metrics
In Section 4.2, we discuss the following metrics: three distance metrics, the Root Mean Square Error (RMSE), the mean bias (BIAS), the coefficient of determination (R2), and one correlation metric, the Pearson Correlation Coefficient (PCC). All the metrics are calculated with the xskillscore library Bell et al. (2021). Definitions are the following:
@@ Appendix B Calculation of verification metrics
Appendix C On the contribution of the VAE
Figure 13 compares power spectra from the COSMO-CLM test reference truth data with power spectra from high-resolution test data generated using three models: the quadratic interpolation of ERA5 data, VAE_res, and LDM_res. This figure highlights the decompression stage’s contribution from the latent space to the pixel space in LDM_res.
VAE_res takes as input the original reference truth high-resolution target variables. LDM_res takes as input the low-resolution ERA5 predictors, the high-resolution static data, and random noise, produces information in the latent space, and projects them in the pixel space using the VAE Decoder. Both VAE_res and LDM_res also use the corresponding UNET estimates for each target variable, but these quantities are subtracted before the encoding step (when applied) and added back after the decompression stage, acting essentially as constants. Therefore, the reconstruction errors in the power spectra for VAE_res are thus solely attributed to the compression/decompression processes, while the reconstruction errors in the power spectra for LDM_res arise from both the diffusion and decompression processes.
For 2-m temperature, the power spectra from the VAE_res and LDM_res are nearly identical across all wavelengths, including the smallest scales. This indicates that the errors in reconstructing the spectra with LDM_res are attributable solely to the decompression stage, with the diffusion process effectively and accurately extracting latent features from the conditional data. On the contrary, for 10-m wind speed, a more chaotic field, discrepancies between the power spectra from LDM_res and VAE_res are observed at scales smaller than approximately 7 km. These differences highlight the errors introduced by the sole extraction of features by the diffusion process.
Figure 13: Comparison of Radially Averaged Power Spectral Density (RAPSD) distributions for results from LDM_res and VAE_res against COSMO-CLM reference-truth and ERA5 quadratic interpolation. The left column refers to the 2-m temperature, and the right column refers to the 10-m wind speed. The first row shows the averaged-in-time spectra, across the whole test dataset. Notice that in the first row y-axes are logarithmic to highlight the tail of the distributions, hence the high frequencies. The bottom rows show the distributions of single-time RAPSD values for fixed wavelengths, namely 269, 20, 9, and 5 km.
Appendix D LDM residual versus non-residual results
In this section, we report the comparison of results from the LDM trained with and without the residual approach, to provide an overview of the improvements introduced by this approach. We compare results in terms of frequency distribution and of Radially Averaged Power Spectral Density, shown in Figures 14 and 15.
@@ Appendix D LDM residual versus non-residual results
The corresponding VAEs for the two models (VAE and VAE_res) perform the same except for the small scales of the 2-m temperature power spectra (not shown). Thus, the lack of performance of LDM can be attributed to the VAE only for this occasion (i.e., (iii)). All the other listed deficiencies are to be solely attributed to the diffusion process itself, which, when trained to reconstruct a residual field instead of the original target variable field, gains performance in the reconstruction of all the frequency distributions, and of the power spectra across all wavelengths (not only the small scales), especially for chaotic variables such as wind speed.
@@
[…]
All authors conceived and conceptualized the study. ET and GF designed the LDM_res architecture and implemented the code for all the tested models. ET managed the data, ran the experiments, performed the analysis and verification of the results, and wrote the manuscript. All authors revised the results and reviewed the manuscript. MC supervised the study from end to end.
[…]
The authors declare that they have no conflict of interest.
Acknowledgements.